# An Information-Theoretic Framework for Receiver Quantization in Communication

### v1

@@ IV-C Geometry of Optimal Uniform Quantization at Receiver
$\mathrm{E}\left[|\mathsf{W}|^{2}\right]=\mathsf{MMSE}=\mathsf{mmse}\cdot\mathrm{E}\left[|\mathsf{V}|^{2}\right]=2\cdot\mathsf{mmse}\cdot\sigma_{v}^{2}.$ (71)
@@
Now we are ready to illustrate the geometry of optimal uniform quantization at the receiver in $N$-dimensional Euclidean space, as shown in Fig. 7. We use boldface letters to denote codewords or signal vectors in the equivalent model (70), e.g., $\mathbf{X}=[\mathsf{X}_{1},...,\mathsf{X}_{N}]$. Since i.i.d. codebook is considered, the vector $\mathbf{X}$ and other vectors in Fig. 7 are all i.i.d. random vectors. Thus, we have $\mathrm{E}\left[\|\mathbf{X}\|^{2}\right]=N\mathcal{E}_{x}$, and the empirical average power of the input codeword converges in probability to $\mathcal{E}_{x}$, i.e.,
@@
Thus, the lengths of $\mathbf{V}$, $\mathbf{W}$, and $\mathbf{Y}$ are $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2})}$, $\sqrt{N\cdot\mathsf{MMSE}}$, and $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2}-\mathsf{MMSE})}$, respectively. We then have two right triangles in Fig. 7: the quantization triangle $\mathbf{0VY}$, which determines the saturation rate $\bar{I}^{*}_{\textrm{GMI}}$ in (66), and the noise triangle $\mathbf{0XV}$, which determines the channel capacity $C$, i.e., the limit of achievable rate under fine quantization. These two triangles jointly determine the triangle $\mathbf{0XY}$, and consequently determine the achievable rate $I^{*}_{\textrm{GMI}}$.
@@
In Fig. 7 there are also some perpendicular lines (dashed) which indicate some projections. For example, the projection of $\mathbf{X}$ onto $\mathbf{V}$ is the linear minimum MSE (LMMSE) estimator of $\mathbf{X}$ from $\mathbf{V}$ as
@@
The channel (84) is a scaling of the channel $Y=\alpha X+D$ in the proof of Proposition 1. We can apply Proposition 1 directly to (84) and obtain exactly the same GMI expression in Theorem 8. Correspondingly, the scaling factor for the nearest neighbor decoder in Fig. 7 should be set as $1-\mathsf{mmse}$.
@@ V Asymptotic Analysis: Characterization of $I_{\textrm{GMI}}(L)$
1) Overload Region ($L<L^{*}$): In this region overload distortion dominates the performance, but the impact of granular distortion is also important when $L$ is close to $L^{*}$. The achievable rate behaves like a waterfall except except for the case where $L$ is very close to $L^{*}$. For high-resolution quantization at high SNR, we have $I_{\textrm{GMI}}\approx\log(L^{3}/4\phi(L))$, which increases approximately linearly with $L$. On the other hand, as $L\to 0$, the effective resolution of the quantizer reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ given in (29).
2) Underload Region ($L>L^{*}$): In this region granular distortion dominates the performance. As $L$ increases, the rate loss increases like $\log(1+L^{2}K^{-2}\mathsf{SNR}/12)$ over a wide range. When $L\gg L^{*}$, the effective resolution reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ again.
@@ V-A Decay of Rate Loss under High-Resolution Receiver Quantization
We first consider the impact of bias in gain control when the irreducible loss is negligible. In other words, we consider how $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{*}\approx C$ as $L\to L^{*}$. Our method is to characterize asymptotic behaviors of rate loss under finite loading factor and finite step size, respectively, in the high-resolution limit. Then the accuracy of the obtained asymptotic formulas is evaluated numerically for finite-resolution receiver quantization.
@@
[…]
@@
is the infimum of the normalized overload distortion $\mathsf{mse}_{\textrm{o}}$ given in (49).
@@
On the other hand, the rate loss due to only granular distortion decays quadratically as the step size vanishes. To prove this we need the following lemma which is one of the various forms of the Euler-Maclaurin summation formula [60].
Lemma 13: For a real-valued continuously differentiable function $f(t)$ defined on $[a,b]$, we have
@@
[…]
@@
is the supremum of the normalized granular distortion $\mathsf{mse}_{\textrm{g}}$ given in (50).
@@
=\;= $\sum\limits_{k=0}^{K-1}\ell\cdot\exp\frac{-k^{2}\ell^{2}}{2}-\frac{\ell}{2}+\frac{\ell}{2}\exp\frac{-K^{2}\ell^{2}}{2}+\frac{K\ell^{3}}{12}\exp\frac{-K^{2}\ell^{2}}{2}+o(\ell^{2})$ (108b)
=\;= $\mathcal{A}+o(\ell^{2}),$ (108c)
@@
=\;= $\sum\limits_{k=0}^{K-1}2k\ell^{2}Q(k\ell)+\frac{\ell^{2}}{8}-\frac{\ell^{2}}{24}+o(\ell^{2})$ (109b)
=\;= $\frac{\mathcal{B}}{\pi}-\frac{\ell^{2}}{24}+o(\ell^{2}).$ (109c)
@@
[…]
Clearly, these approximations successfully capture the decay of rate loss when it is dominated by the overload distortion. When the loading factor $L$ exceeds $4$, the increasing granular distortion kicks in and dominates the performance quickly. In this case the achievable rate is well approximated by $L^{2}\mathsf{SNR}/12K^{2}$. When the resolution decreases, e.g., in Fig. 9 where the resolution is 6 bits, the approximation (112) becomes less useful, especially at high SNR. But the approximation (113) is still satisfactory. For higher accuracy, by considering both the overload distortion and the granular distortion, we propose an approximation given by
@@ V-B Properties and Approximations of $I_{\textrm{GMI}}(L)$
We first discuss some general properties of $I_{\textrm{GMI}}(L)$. The derivative of the achievable rate
@@
which does not varies with the SNR, implying that the achievable rate saturates; see high-SNR curves in Fig. 4(e) and Fig. 4(f).
@@ V-B2 Underload Region
$I_{\textrm{GMI}}(L)$ $\approx C-\log\left(1+\frac{12L^{2}}{K^{2}}\mathsf{SNR}\right),$ (121)
[…]
$\bar{I}_{\textrm{GMI}}(L)=\log(1/\gamma)\approx\log\frac{K^{2}}{12L^{2}}.$ (122)
@@ VI-A Approximations of Optimal Loading Factor
$\lim\limits_{K\to\infty}\frac{24K^{2}\exp\frac{{-L^{\ast}}^{2}}{2}}{\sqrt{2\pi}{L^{\ast}}^{3}}=1.$ (126)
@@
Remark: In fact, the equation (135) can also be derived from analysis of the MSE: Just let the derivative of the MSE approximation (115) be zero, and omit the exponential terms therein except for the dominated one.
@@
In Fig. 12, we show $L^{\ast}$, its approximation $\hat{L}_{0}$ from (131) in Proposition 15, the simple approximation $2\sqrt{\ln(2K)}$, and the approximations $\hat{L}_{2}$ and $\hat{L}_{3}$ from [46]. Note that the case $b=1$ does not require gain control. We see that the new approximation $\hat{L}_{0}$ is accurate for all resolutions $b\geq 2$. It is much better than $\hat{L}_{2}$ and close to the more complicated approximation $\hat{L}_{3}$.
Remark: To approach $I^{*}_{\textrm{GMI}}$, it is sufficient to use a simple linear approximation of $L^{*}$ as
@@ VI-B Maximum Achievable Rate Approximations and Per-Bit Rules
We have found a simple relationship between $I_{\textrm{GMI}}^{*}$ and $\mathsf{mmse}$ in Theorem 8, and we have also given numerical results of $\mathsf{mmse}$ for $b=1,2,...,12$ in Table I. However, a direct connection between $I_{\textrm{GMI}}^{*}$ and the resolution is still very useful. The following result provides such a connection via approximating $\mathsf{mmse}$ by $b$.
@@
$\lim\limits_{K\to\infty}\frac{3K^{2}}{\ln(2K)}\mathsf{mmse}=1,$ (139)
which yields a high-resolution approximation of MMSE given by
@@
The approximation (140) is still not accurate enough for moderate to low resolutions; see Table I. More approximations for MMSE or $\mathsf{SNR}_{q}$ can be found in [References, Sec. IV] based on refined asymptotic formulas of $L^{*}$ and its asymptotic relationship with MMSE. However, using (140) is enough to get the simple and useful approximation of $I_{\textrm{GMI}}^{*}$ in Proposition 16. In Fig. 14 we compare $I^{*}_{\textrm{GMI}}$ with its approximation $\hat{I}_{\textrm{GMI}}$. It is shown that the accuracy is acceptable when $b\geq 2$.
@@
[…]
Numerical evaluations of $\hat{\bar{I}}_{\textrm{GMI}}(b)$ for $1\leq b\leq 12$ are shown in Table I. In Fig. 14, it is shown that $\mathsf{mmse}$ decreases exponentially as the resolution increases, and correspondingly, the saturation rate $\bar{I}^{*}_{\text{GMI}}$ grows approximately linearly with the resolution. The improvement per bit is $1.6$-$1.9$ bits/c.u. in our range of interest, although the high-resolution limit is $2$ bits/c.u.
Figure 13: Approximation of $I^{*}_{\textrm{GMI}}$.
@@ VII-A Relationship Between MSE and Achievable Rate: Numerical Examples
[…]
@@
For the channel (4) with a uniform quantizer $q(\cdot)$ whose resolution is at least two bits, we have shown that the unique gain control factor that minimizes the MSE also maximizes the GMI. When non-uniform quantization rules are allowed, one may vary the thresholds $\{\ell_{k}\}$ and levels $\{y_{k}\}$ of $q(\cdot)$ (possibly under some constraints) to reduce $\gamma$ so that the GMI (19) can be optimized, yielding thresholds $\{\ell_{k}^{*}\}$ and levels $\{y^{*}_{k}\}$. However, when this optimized quantization rule is applied in a given channel, to achieve that optimized GMI, we need to set the gain control factor $g$ appropriately according to $\sigma_{v}$ (the standard deviation of the quantizer input) to guarantee that the thresholds satisfy $\ell_{k}=l_{k}/g=\ell^{*}_{k}$ for $1\leq k<K$. Otherwise, the performance will be degraded. We next explore the impact of gain control on the MSE and the GMI of the channel (4) under nonuniform quantization rules by examples; see numerical results given in Fig. 17, where we set $\mathsf{SNR}=10$ dB. The examples are described as follows, each of which satisfies $42K=$ 4, i.e., $b=2$.
@@
From Fig. 17(a), we confirm that, under uniform quantization, the parameter $\gamma$ converges to $\mathsf{mmse}$ when the loading factor approaches its optimal value $L^{*}$ (in this example it is $2.3441$); see also Fig. 17(d). Interestingly, this consistency also occurs in Fig. 17(b), where we use the optimal $8$-level quantizer obtained numerically in [30] by Lloyd’s algorithm [70]. But for the nonuniform quantizer given in Fig. 17(c) and the optimized quantizer with equispaced thresholds (also obtained in [30]) given in Fig. 17(e), the gain control factor that maximizes the GMI and the one that minimizes the MSE are different. The highly nonlinear quantizer given in Fig. 17(f) provides an example of unusual behavior of $I_{\textrm{GMI}}(g)$, which achieves its maximum when the MSE is relatively large.
@@
$\hat{R}$ $=\log\left(1+\frac{\mathcal{E}_{x}}{\sigma^{2}+2\cdot\mathsf{mse}\cdot\sigma_{v}^{2}}\right)$ (148)
$=\log(1+\mathsf{SNR})-\log\left(1+\frac{\mathsf{mse}}{1+\mathsf{mse}}\mathsf{SNR}\right),$ (149)

### v2

@@ IV-C Geometry of Optimal Uniform Quantization at the Receiver
[…]
@@
Now we are ready to illustrate the geometry of optimal uniform quantization at the receiver in the $N$-dimensional Euclidean space, as shown in Fig. 7. We use boldface letters to denote codewords or signal vectors in the equivalent model (69), e.g., $\mathbf{X}=[\mathsf{X}_{1},...,\mathsf{X}_{N}]$. Since i.i.d. codebook is considered, the vector $\mathbf{X}$ and other vectors in Fig. 7 are all i.i.d. random vectors. Thus, we have $\mathrm{E}\left[\|\mathbf{X}\|^{2}\right]=N\mathcal{E}_{x}$, and the empirical average power of the input codeword converges in probability to $\mathcal{E}_{x}$, i.e.,
@@
Thus, the lengths of $\mathbf{V}$, $\mathbf{W}$, and $\mathbf{Y}$ are $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2})}$, $\sqrt{N\cdot\mathsf{MMSE}}$, and $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2}-\mathsf{MMSE})}$, respectively. We then have two right triangles in Fig. 7: the quantization triangle $\mathbf{0VY}$ [^9] Note that the angle between $\mathbf{Y}$ and $\mathbf{V}-\mathbf{Y}$ is a right angle. Since Fig. 7 is a one-dimensional illustration of the geometry in the $N$-dimensional space, we cannot make every right angle therein look like a right angle., which determines the saturation rate $\bar{I}^{*}_{\textrm{GMI}}$ in (65), and the noise triangle $\mathbf{0XV}$, which determines the channel capacity $C$, i.e., the limit of achievable rate under fine quantization. These two triangles jointly determine the triangle $\mathbf{0XY}$, and consequently determine the achievable rate $I^{*}_{\textrm{GMI}}$.
@@
In Fig. 7 there are also some perpendicular lines (dashed) which indicate some projections. For example, the projection of $\mathbf{X}$ onto $\mathbf{V}$ is the LMMSE estimator of $\mathbf{X}$ from $\mathbf{V}$ as
@@
The channel (83) is a scaling of the channel $Y=\alpha X+D$ in the proof of Proposition 1. We can apply Proposition 1 directly to (83) and obtain exactly the same GMI expression as (54) in Theorem 8. Correspondingly, the scaling factor for the nearest neighbor decoder in Fig. 7 should be set as $1-\mathsf{mmse}$ to achieve (54), while in the original model (Fig. 1) we should set it according to (16).
@@ V Asymptotic Analysis: Characterization of $I_{\textrm{GMI}}(L)$
1) Overload Region ($L<L^{*}$): In this region overload distortion dominates the performance and the achievable rate behaves like a waterfall. In particular, for high-resolution quantization at high SNR, the GMI is approximated by $\log(L^{3}/4\phi(L))$ (except when $L$ is very close to $L^{*}$), which appears almost linear; see (119a) and Fig. 5, (d)-(f). When $L$ approaches $L^{*}$, the impact of granular distortion becomes increasingly important, so that the GMI grows slower and approaches its maximum. As $L\to 0$, the effective resolution of the quantizer reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ given in (35).
2) Underload Region ($L>L^{*}$): In this region granular distortion dominates the performance. As $L$ increases, the rate loss increases like $\log(1+L^{2}K^{-2}\mathsf{SNR}/12)$ (see (120)) over a wide range. When $L\gg L^{*}$, the effective resolution reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ again.
@@ V-A Decay of Rate Loss under High-Resolution Receiver Quantization
Now consider the impact of bias in gain control when the irreducible loss is negligible. In other words, we consider how $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{*}\approx C$ as $L\to L^{*}$. Our method is to characterize asymptotic behaviors of rate loss under finite loading factor and finite step size, respectively, in the high-resolution limit. Then the accuracy of the obtained asymptotic formulas is evaluated numerically for finite-resolution receiver quantization.
@@
$=\left((1+o_{L}(1))\frac{4\phi(L)}{L^{3}}\right)\mathsf{SNR}\;\;\textrm{nats/c.u.}$ (94b)
@@
is the infimum of the normalized overload distortion $\mathsf{mse}_{\textrm{o}}$ given in (48a).
@@
On the other hand, the rate loss due to only granular distortion decays quadratically as the step size vanishes. To prove this we need the following lemma which is one of the various forms of the Euler-Maclaurin summation formula.
Lemma 13[61]: For a real-valued continuously differentiable function $f(t)$ defined on $[a,b]$, we have
@@
$C-I_{\text{GMI}}$ $=\left(\overline{\mathsf{mse}}_{\textrm{g}}+o(\ell^{2})\right)\mathsf{SNR}\;\;\textrm{nats/c.u.}$ (105a)
$=\left(\frac{\ell^{2}}{12}+o(\ell^{2})\right)\mathsf{SNR}\;\;\textrm{nats/c.u.}$ (105b)
@@
is the supremum of the normalized granular distortion $\mathsf{mse}_{\textrm{g}}$ given in (49a).
@@
$=\;$ $\sum\limits_{k=0}^{K-1}\ell\cdot\exp\frac{-k^{2}\ell^{2}}{2}-\frac{\ell}{2}+\frac{\ell}{2}\exp\frac{-K^{2}\ell^{2}}{2}+\frac{K\ell^{3}}{12}\exp\frac{-K^{2}\ell^{2}}{2}+o(\ell^{2})$ (107b)
$=\;$ $\mathcal{A}+o(\ell^{2}),$ (107c)
@@
$=\;$ $\sum\limits_{k=0}^{K-1}2k\ell^{2}Q(k\ell)+\frac{\ell^{2}}{8}-\frac{\ell^{2}}{24}+o(\ell^{2})$ (108b)
$=\;$ $\frac{\mathcal{B}}{\pi}-\frac{\ell^{2}}{24}+o(\ell^{2}).$ (108c)
@@
$C-I_{\textrm{GMI}}(L)\approx L^{2}\mathsf{SNR}/12K^{2}\;\;\textrm{nats/c.u.}$ (112)
Clearly, these approximations successfully capture the decay of rate loss when it is dominated by the overload distortion. When the loading factor $L$ exceeds $4$, the increasing granular distortion kicks in and dominates the performance quickly and the rate loss is well approximated by (112). When the resolution decreases, e.g., in Fig. 9 where the resolution is 6 bits, the approximation (111) becomes less useful, especially at high SNR. But the approximation (112) is still satisfactory. For higher accuracy, by considering both the overload distortion and the granular distortion, we propose an approximation given by
@@ V-B Properties and Approximations of $I_{\textrm{GMI}}(L)$
Now discuss some general properties of $I_{\textrm{GMI}}(L)$. The derivative of the achievable rate
@@
which does not vary with the SNR, implying that the achievable rate saturates; see high-SNR curves in Fig. 4(e) and Fig. 4(f).
@@ V-B2 Underload Region
$I_{\textrm{GMI}}(L)$ $\approx C-\log\left(1+\frac{L^{2}}{12K^{2}}\mathsf{SNR}\right),$ (120)
which is highly accurate even when the resolution is low; see Fig. 11 and Fig. 11. We have also considered an even simpler approximation $C-L^{2}K^{-2}\mathsf{SNR}/12\;\textrm{nats/c.u.}$ in the high-resolution scenario; see Fig. 9 and Fig. 9. However, it becomes less useful as the loading loss increases, as shown in Fig. 11 and Fig. 11. The approximation (120) also implies that the saturation rate can be approximated by
$\bar{I}_{\textrm{GMI}}(L)=\log(1/\gamma)\approx\log\frac{12K^{2}}{L^{2}}.$ (121)
@@ VI-A Approximations of Optimal Loading Factor
$\lim\limits_{K\to\infty}\frac{24K^{2}\exp\frac{{-L^{\ast}}^{2}}{2}}{\sqrt{2\pi}{L^{\ast}}^{3}}=1,$ (125)
which is equivalent to
$\lim\limits_{K\to\infty}\frac{2\ln(2K)+\ln 6}{{L^{\ast}}^{2}/2+3\ln{L^{\ast}}+\ln\sqrt{2\pi}}=1.$ (126)
@@
Remark: The equation (132) coincides with (126) in the aforementioned new proof of (61). In fact, (132) or (135) can also be heuristically argued via an approximate analysis of the MSE: Just let the derivative of the MSE approximation (114) be zero, and omit the exponential terms therein except for the dominated one.
@@
In Fig. 12, we show $L^{*}$, its approximation $\hat{L}_{0}$ from (131) in Proposition 15, the simple approximation $2\sqrt{\ln(2K)}$, and the approximations $\hat{L}_{2}$ and $\hat{L}_{3}$ from [48]. Note that the case $b=1$ does not require gain control. We see that the new approximation $\hat{L}_{0}$ is accurate for all resolutions $b\geq 2$. It is much better than $\hat{L}_{2}$ and close to the more complicated approximation $\hat{L}_{3}$. Interestingly, for moderate resolutions it can be observed that $L^{*}$ increases approximately linearly. Therefore, to approach $I^{*}_{\textrm{GMI}}$, it is sufficient to use a simple linear approximation of $L^{*}$ as
@@ VI-B Maximum Achievable Rate Approximations and Per-Bit Rules
We have found a simple relationship between $I_{\textrm{GMI}}^{*}$ and $\mathsf{mmse}$ in Theorem 8, and we have also given numerical results of $\mathsf{mmse}$ for b=1,2,,12b=1,2,...,12 in Table I. However, a direct connection between $I_{\textrm{GMI}}^{*}$ and the resolution is still very useful. The following result provides such a connection via approximating $\mathsf{mmse}$ by $b$.
@@
$\lim\limits_{K\to\infty}\frac{3K^{2}}{\ln(2K)}\mathsf{mmse}=1.$ (139)
Replacing $K$ by $2^{b-1}$ yields a high-resolution approximation of MMSE given by
@@
The approximation (140) is still not accurate enough for moderate to low resolutions; see Table I. More approximations for MMSE or $\mathsf{SNR}_{q}$ can be found in [References, Sec. IV] based on refined asymptotic formulas of $L^{*}$ and its asymptotic relationship with MMSE. However, using (140) is enough to get the simple and useful approximation of $I_{\textrm{GMI}}^{*}$ in Proposition 16. In Fig. 13 we compare $I^{*}_{\textrm{GMI}}$ with its approximation $\hat{I}_{\textrm{GMI}}(b)$. It is shown that the accuracy is acceptable when $b\geq 2$.
Figure 13: Approximation of $I^{*}_{\textrm{GMI}}$.
@@
    $\lim\limits_{b\to\infty}\left(\bar{I}_{\textrm{GMI}}^{*}(b)-\bar{I}_{\textrm{GMI}}^{*}(b-1)\right)=2\;\textrm{bits/c.u.}$ (145)
Numerical evaluations of $\hat{\bar{I}}_{\textrm{GMI}}(b)$ for $1\leq b\leq 12$ are shown in Table I. In Fig. 14, it is shown that $\mathsf{mmse}$ decreases exponentially as the resolution increases, and correspondingly, the saturation rate $\bar{I}^{*}_{\text{GMI}}$ grows approximately linearly with the resolution. The improvement per bit is 1.61.6-$1.9$ bits/c.u. in our range of interest, although the high-resolution limit is $2$ bits/c.u.
@@ VII-A Relationship Between MSE and Achievable Rate: Numerical Examples
(a) Uniform quantizer:              thresholds $0,\pm 0.25,\pm 0.5,\pm 0.75,\mspace${108.0mu} levels $\pm 0.125,\pm 0.375,\pm 0.625,\pm 0.875$
(b) Optimal nonuniform quantizer:        thresholds $0,\pm 0.5006,\pm 1.0500,\pm 1.7480,\mspace${48.0mu} levels $\pm 0.2451,\pm 0.7560,\pm 1.3440,\pm 2.1520$
(c) Nonuniform quantizer:            thresholds $0,\pm 0.2,\pm 0.6,\pm 0.7,\mspace${128.0mu} levels $\pm 0.1,\pm 0.5,\pm 0.7,\pm 0.9$
(d) Part of (a): $2.1\leq L\leq 2.6\mspace${128.0mu}
(e) Optimized quantizer with equispaced thresholds:                 thresholds $0,\pm 0.5646,\pm 1.1292,\pm 1.6937,\mspace${48.0mu} levels $\pm 0.2749,\pm 0.8247,\pm 1.3747,\pm 2.1049$
(f) Nonuniform quantizer:            thresholds $0,\pm 0.2,\pm 0.6,\pm 0.7,\mspace${128.0mu} levels $\pm 0.1,\pm 0.9,\pm 0.8,\pm 0.2$
@@
For the channel (4) with a uniform quantizer $q(\cdot)$ whose resolution is at least two bits, we have shown that the unique gain control factor that minimizes the MSE also maximizes the GMI. When non-uniform quantization rules are allowed, one may vary the thresholds $\{\ell_{k}\}$ and levels $\{y_{k}\}$ of $q(\cdot)$ (possibly under some constraints) to reduce $\gamma$ so that the GMI (21) can be optimized, yielding thresholds $\{\ell_{k}^{*}\}$ and levels $\{y^{*}_{k}\}$. However, when this optimized quantization rule is applied in a given channel, to achieve that optimized GMI, we need to set the gain control factor $g$ appropriately according to $\sigma_{v}$ (the standard deviation of the quantizer input) to guarantee that the thresholds satisfy $\ell_{k}=l_{k}/g=\ell^{*}_{k}$ for $1\leq k<K$. Otherwise, the performance will be degraded. We next explore the impact of gain control on the MSE and the GMI of the channel (4) under nonuniform quantization rules by examples; see numerical results given in Fig. 17, where we set $\mathsf{SNR}=10$ dB. The examples are described as follows, each of which satisfies $2K=4$, i.e., $b=2$.
@@
From Fig. 17(a), we confirm that, under uniform quantization, the parameter $\gamma$ converges to $\mathsf{mmse}$ when the loading factor approaches its optimal value $L^{*}$ (in this example it is $2.3441$); see also Fig. 17(d). Interestingly, the consistency also occurs in two examples: Fig. 17(b) for the optimal $8$-level quantizer obtained numerically in [31] by Lloyd’s algorithm [71], and Fig. 17(e) for the optimized quantizer with equispaced thresholds (also obtained in [31]). But for the nonuniform quantizer given in Fig. 17(c) and Fig. 17(f), the gain control factor that maximizes the GMI and the one that minimizes the MSE are different. The highly nonlinear quantizer given in Fig. 17(f) provides an example of unusual behavior of $I_{\textrm{GMI}}(g)$, which achieves its maximum when the MSE is relatively large.
@@
$\hat{R}$ $=\log\left(1+\frac{|h|^{2}\sigma_{x}^{2}}{\sigma^{2}+2\cdot\mathsf{mse}\cdot\sigma_{v}^{2}}\right)$ (148a)
$=\log(1+\mathsf{SNR})-\log\left(1+\frac{\mathsf{mse}}{1+\mathsf{mse}}\mathsf{SNR}\right),$ (148b)